Status and progress of installation related activities

This page last update -- GianlucaAglieriRinella - 23 Oct 2008

  • Infrastructure at point 2
    • (TO DO) Request and delivery of the PIT Operator Node computer (Cesar to remind Peter Chochula)
    • Crate in rack C22 installed
      • Power supply repaired and connected (DONE)
      • Cooling connected (DONE)
      • CANbus cable connection to remote control power supply missing (DONE)
    • Optical patch panel in rack C22 (splitters)
      • Installed (DONE)
      • Optical connections from detector to be commissioned (DONE)
    • Optical TTC link (DONE)
      • Fiber rounting from CTP fanout to crate (DONE)
        • WARNING: this may need an optical attenuator, see CTP requirements
    • (DONE) Windows DCS working node (alipitwn00 ?) to control power supply (CAN) and to run PVSS project
      • Installation (DONE)
      • Our PVSS project installation (DONE)
      • CANbus controller and cable connection (DONE) * (DONE) SLC4 DCS working node to host the dim_fedServer software
      • PC (formerly DAQ pc aldaqpc021 in DSF, MAC:00-30-48-2D-71-42) installed in rack in CR3 (DONE by Fuchs)
      • DDL optical link between this PC and crate in rack C22 (DONE)
      • DCS SLC4 installation, renaming of PC to DCS convention (alipitwn01 ?) (DONE)
      • DATE installation, fecc2rorc library (DONE)
      • fedServer installation (DONE)
    • (DONE) Dedicated laptop to temporarily allow remote reconfiguration
      • laptop procurement (DONE)
      • laptop registration (DONE)
      • NICE and SLC4 installation (DONE)
      • Xilinx USB programmer procurement (DONE)
      • Installation of XIlinx programming software (DONE)

  • Production BRAIN board
    • Reparation at CERN of instance 1 (DONE)
    • JTAG access test (DONE)
    • Full JTAG boundary interconnection test (DONE)
    • Clock verification (DONE)
    • Commissioning
      • Fast-OR buses BER measurements (DONE)
      • Control bus communication BER measurement (DONE)
      • Test with detector modules in DSF (DONE)
      • Operation with 10 OPTIN plugged, UDF Fast-OR pattern test, temperature measurement (TO DO)

  • Activity in DSF
    • 4 half staves installed, 2 hs operational (tested with full readout)
    • Test system
      • (DONE)Interlock in place
      • (DONE)spdFed usable version
      • (DONE) SPD PVSS
    • (DONE) Transfer of pixeltrigger setup from pixel lab to DSF
    • (DONE) Testing of the full chain, cosmic run with pixeltrigger as trigger source
    • (DONE) Measurement of pit latency (DONE)
    • (DONE) Operation of 10 OPTIN boards simultaneously plugged, automated BER test
    • (DONE) Tests of the TTCrx reset procedure
      • Program BRAIN V1 and test the reset of TTC
      • Test reset from software

  • Other
    • Procurement of more 12-LC/MTP optical cables, only one spare left after installation at point 2 (TO DO)
    • Measurements of Fast-OR latencies with laser pulse (summer student)
Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r8 - 2008-10-23 - GianlucaAglieriRinella
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    AliceSPD All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback